OpenStudy (anonymous):

A line contains the points (–4, 5) and (3, –9). Write the equation of the line using point-slope form

OpenStudy (p0sitr0n):

\[y=mx+b, m=\frac{ y_2-y_1 }{ x_2-x_1 }\]

OpenStudy (anonymous):

A. Y+2=-2(X-3) B.Y+5=-2(X-4) C.Y-3=-2(X+9) D.Y-5=-2(X+4)

OpenStudy (p0sitr0n):

thus the slope is m=-9-5 / 3- -4=-14/7=-2 y=-2x+b plug in one of the points, say 3,-9 -9=-6+b b=-3 thus, y=-2x-3

OpenStudy (anonymous):

no I need one of those

OpenStudy (anonymous):

is it a

OpenStudy (anonymous):

@P0sitr0n

OpenStudy (anonymous):

\[y-y _{1}=\frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\left( x-x _{1} \right)\]

OpenStudy (anonymous):

dude no

OpenStudy (anonymous):

is a right

OpenStudy (anonymous):

A. Y+2=-2(X-3) B.Y+5=-2(X-4) C.Y-3=-2(X+9) D.Y-5=-2(X+4)

OpenStudy (anonymous):

no

OpenStudy (anonymous):

how

OpenStudy (anonymous):

it has to be y-5 or y+9 on L.H.S

OpenStudy (anonymous):

so b

OpenStudy (anonymous):

or d

OpenStudy (anonymous):

check which can be, not guess.

OpenStudy (anonymous):

im canceling out

OpenStudy (anonymous):

y-y1 means either y-5 or y+9 on L.H.S x-x1 on R.H.S can be x+4 or x-3

OpenStudy (anonymous):

ik yur job isnt to just give me the anwers but im tired and ive been working all day at 8am i was up working on this math stuff and im really tired of looking at math and i have another house and a half of math work so can yu cut me some

OpenStudy (anonymous):

hour*

OpenStudy (anonymous):

m=-2 as Mr. P0sitr0n calculated. y-5=-2(x+4) matches it.

OpenStudy (anonymous):

yu said it wasnt a

OpenStudy (anonymous):

oh nvm thank you soo much

OpenStudy (anonymous):

inA it is y+2